# Not Only Rewards But Also Constraints: Applications on Legged Robot Locomotion

### v2

@@ I Introduction
Figure 1: Several legged robots trained with our learning framework to locomote in challenging terrains
@@
Unlike other related works [10, 11] that focused on safety for the benefit of constraints, our work investigates a new advantage of utilizing constraints from a practical engineering perspective. To the best of our knowledge, our work is the first to show highly robust real-world robot performance for multiple legged robots using constrained reinforcement learning and demonstrate the benefits of constraints from an engineering standpoint.
@@ I-A Contributions
  • Through extensive simulation and real-world experiments involving multiple legged robots with varying morphologies and physical properties, we demonstrate the capability of leveraging constraints in the learning pipeline to improve robot generalizability and engineering intuitiveness, which are difficult to achieve when relying solely on rewards, while maintaining highly robust control performance.
@@
  • We make our implementations publicly available at https://github.com/railabatkaist/legged-robot-constrained-rl for future works on applications on different robot platforms/tasks and algorithmic improvements.
@@ II-B Legged robot locomotion
There are some previous works to reduce the burden of reward engineering. Escontrela et al. [32], Wu et al. [33], and Fuchioka et al. [34] utilized an auxiliary task to imitate the natural motions present in expert demonstration data, derived either from animal motions or trajectory optimization. However, the motions that the controller can generate are limited to the available expert demonstration data. Feng et al. [35] proposed a morphology randomization technique to generate a general controller for quadruped robots and remove the robot-specific engineering process. However, training a single controller for a broad range of quadruped robot morphologies results in a conservative controller.
@@ II-C Constrained reinforcement learning
Our policy optimization algorithm is based on IPO but several improvements are added to make the algorithm scalable for multiple constraints. First, an adaptive constraint thresholding method is used to appropriately set the constraint limits based on the current policy’s performance and guide the policy to the constraint-satisfying region using steep gradients of logarithmic barrier functions. Second, a multi-head cost value function is proposed to parallelize cost value function training and cost advantage computation in a memory-efficient manner. With these improvements, we demonstrate that our framework is capable of training controllers for a variety of legged robots, which necessitates the incorporation of more than ten constraints. Furthermore, we showcase the robustness of these controllers in real-world scenarios. Our strong experimental results demonstrate that our method is superior for real-world robotic applications when compared to the prior works [42], [43] that were tested on simulated benchmark tasks [47] with one or two constraints. There is a previous work by Gangapurwala et al. [10] that defined the problem of training locomotion controllers for legged robots in a CMDP formulation. However, their framework required cost coefficient engineering because they handled several constraints as rewards with appropriately set coefficients. Furthermore, their learning framework required considerable reward engineering due to multiple reward terms and a reference trajectory generator, whereas our framework requires only three reward terms and no demonstration data.
@@ IV-B Policy optimization
In the original paper [43], the IPO objective (10) is optimized with Proximal Policy Optimization (PPO) [49]. However, we utilized Trust Region Policy Optimization (TRPO) [48] because the optimizer’s explicit line search allows for more stable policy improvement and constraint feasibility checking  [^1] We included the version using the PPO [49] optimizer in the publicly available implementation of our method so that users can benefit from the algorithm’s advantages, such as the ease of implementation and the availability of recurrent neural networks.. The advantage functions $A^{\pi}$ and $A^{\pi}_{C_{k}}$ are computed using Generalized Advantage Estimation (GAE) [50] and neural network value functions. Each advantage function then undergoes additional standardization and zero-mean normalization before computing the policy gradients.
@@ VI Experimental Results
  • $*{}^{*}$esults are after training with 1.6 billion time steps. Constraint thresholds are reported in the top row or inside the bracket. Bolded texts indicate violated or barely satisfied constraints.
@@ VI-A Implementation detail
The locomotion control policy runs at 100Hz. We deployed the trained controller on two real-world quadruped robots: Raibo [9] and Mini-cheetah [57]. For the deployment, the student policy network was reimplemented with the Eigen library in C++ and forward-passed in the Central Processing Unit (CPU). We recommend readers check the original paper [9, 57] for more details about the robot hardware. Hyperparameters are provided in the Appendix (Table IX, X).
@@ VI-B Framework Evaluation
The trained controllers were deployed on two real-world quadruped robots available in the lab: Raibo and Mini-cheetah. The controllers were extensively tested in various harsh terrains both indoor and outdoor to verify the robust control performance as shown in the previous work [6]. The testing environments included slippery hills, steep slopes, deformable terrains, stairs, a pile of leaves with hidden impediments, a moving cart, unstably piled planks, and even more. No matter how difficult the terrains were, Raibo and Mini-cheetah were able to walk steadily by modifying base motions and foot placements based solely on proprioceptive sensor data, as illustrated in Figure 4. Furthermore, Raibo could sprint over grassy terrain at 2m/s and climb slippery and steep slopes with an angle of around ${}^{\circ}$ at 1m/s which is roughly double the speed reported in the prior work [6].
@@ VI-D Comparison with the Reward-only Framework
We compare the generalizability of our framework and the reward-only framework, which is dominantly used in previous works on legged robot locomotion [6, 21, 8, 9]. The reward terms and hyperparameters used for the reward-only framework are based on the previous works by Ji et al.[8] and Choi et al. [9], and are laboriously engineered for robust control performance on the Raibo robot. There were a total of twelve reward terms utilized, each with highly calibrated reward coefficients. The reward terms are command tracking, air time, foot clearance, orthogonal velocity, torque usage, joint velocity, joint acceleration, nominal configuration, hip joint usage, action smoothness, foot slippage, and undesired body contact. TRPO was used for policy optimization.
@@
Figure 7: Comparison of each reward component’s distribution with the reward-only framework. Red blocks show reward types with extremely different reward distributions.
@@ VI-E2 Multi-head cost value function
  • $*{}^{*}$ Percentage in the ”Computation time” indicates the ratio of time spent on cost-critic versus others (e.g., simulation), where others are identical for all three methods.
The results are shown in Figure 8-B and Table VI. The multi-head method requires fewer trainable parameters than the single-head method. The multi-head cost value function requires 0.1 million trainable parameters, whereas single-head cost value functions require 1 million. Despite having considerably fewer neural network parameters to optimize, the policy optimized with the multi-head version yielded a similar reward as the single-head version (Figure 8-B). This indicates that sharing the neural network backbone is permissible for each cost-value function, and the network capacity is sufficient to model them. Because only the output layer size increases in proportion to the number of constraints, the suggested technique is extremely scalable and effective for solving problems with numerous constraints.
@@ VII-A Comparison of velocity command tracking performance with the reward-only framework
As shown in Table VII, we can achieve the same command tracking performance with similar torque usage within our framework by applying the same exponential function for the command tracking reward and adjusting the reward coefficients accordingly ($k_{c}=200$, $\hat{k}_{\tau}=0.03$). When compared to a quadratic function, an exponential function has a steeper gradient near zero, resulting in more sensitive reward feedback for different command tracking errors. As a result, the robot can be trained to exhibit smaller command tracking errors.
@@
Ours + L2 0.498 ($\pm$ 0.308) 24.769 ($\pm$ 5.811)
Ours + Exp 0.367 ($\pm$ 0.290) 29.489 ($\pm$ 8.311)
Reward-only 0.359 ($\pm$ 0.354) 30.773 ($\pm$ 10.181)
@@ VII-B Extra demonstration of constraint’s engineering advantages
  • $*{}^{*}$ ”(A)” indicates the engineer’s action. ”(R)” indicates the engineer’s response after reviewing the trained results for the action taken in ”(A)”.
@@
VII-C Hyperparameters
The hyperparameters used for the experiments are specified below.
TABLE IX: Environment hyperparameters
Number of environments 200
Episode length 4 $\mathrm{s}$
Terrain Hills Frequency [0.2, 1.0]
Amplitude [0.2, 1.4]
Discrete hills Amplitude [0.2, 1.2]
Step size [0.02, 0.15]
Steps Width [0.1, 0.5]
Height [0.05, 0.2]
Inclining steps Roughness [0.05, 0.2]
Height [0.02, 0.1]
Stairs Width [0.2, 0.5]
Height [0.02, 0.16]
PD gains Raibo (50.0, 0.5)
Mini-cheetah (17.0, 0.4)
Hound (50.0, 1.0)
Anymal B (50.0, 0.2)
Anymal C (50.0, 0.2)
Go1 (20.0, 0.5)
Atlas (300.0, 3.0)
*(200.0, 2.0) for ankle joints
TABLE X: Teacher-Student learning hyperparameters
Discount factor 0.99
Maximum gradient norm 1
Value epochs 20
Learning rate 3e-4
GAE coefficient 0.97
Entropy coefficient 0.05
Policy optimizer Maximum step size 0.01
Line search decay rate 0.8
Damping coefficient 0.01
Conjugate gradient iterations 10
Teacher network MLP encoder [80, 60, *24]
MLP actor [256, 160, 128, *12]
Standard deviation 12
Student network GRU encoder [128, *24]
MLP actor [256, 160, 128, *12]
Value network [160, 128, *1]
Multi-head cost value network [160, 128, *10]
Activation function leaky-relu
  • $*{}^{*}$ The last values in the network parameters are the output dimensions.

### v3

@@ I Introduction
Figure 1: Several legged robots trained with our learning framework to locomote in challenging terrains. The robots include six quadrupedal robots and one bipedal robot.
@@
Unlike other related works [10, 11] that focused on safety for the benefit of constraints, our work investigates a new advantage of utilizing constraints from a practical engineering perspective. To the best of our knowledge, our work is the first to show highly robust simulation and real-world robot performance for multiple legged robots (i.e., six quadrupedal robots and one bipedal robot) using constrained reinforcement learning and demonstrate the benefits of constraints from an engineering standpoint.
@@ I-A Contributions
  • Through extensive simulation and real-world experiments involving multiple legged robots with varying morphologies and physical properties (i.e., six quadrupedal robots and one bipedal robot), we demonstrate the capability of leveraging constraints in the learning pipeline to improve robot generalizability and engineering intuitiveness, which are difficult to achieve when relying solely on rewards, while maintaining highly robust control performance.
@@ II-B Legged robot locomotion
There are some previous works to reduce the burden of reward engineering. Escontrela et al. [32], Wu et al. [33], and Fuchioka et al. [34] utilized an auxiliary task to imitate the natural motions present in expert demonstration data, derived either from animal motions or trajectory optimization. However, the motions that the controller can generate are limited to the available expert demonstration data. Feng et al. [35] proposed a morphology randomization technique to generate a general controller for quadruped robots and remove the robot-specific engineering process. However, training a single controller for a broad range of quadruped robot morphologies results in a conservative controller and is a well-known problem of domain randomization techniques. Furthermore, their experiments were restricted to indoor settings with simple behaviors (e.g., go forward, turn in place), limiting their ability to demonstrate high controller robustness.
@@ II-C Constrained reinforcement learning
Our policy optimization algorithm is based on IPO but several improvements are added to make the algorithm scalable for multiple constraints. First, an adaptive constraint thresholding method is used to appropriately set the constraint limits based on the current policy’s performance and guide the policy to the constraint-satisfying region using steep gradients of logarithmic barrier functions. Second, a multi-head cost value function is proposed to parallelize cost value function training and cost advantage computation in a memory-efficient manner. With these improvements, we demonstrate that our framework is capable of training controllers for a variety of legged robots, which necessitates the incorporation of more than ten constraints. Furthermore, we showcase the robustness of these controllers in real-world scenarios. Our strong experimental results demonstrate that our method is superior for real-world robotic applications when compared to the prior works [42], [43] that were tested on simulated benchmark tasks [47] with one or two constraints. There is a previous work by Gangapurwala et al. [10] that defined the problem of training locomotion controllers for legged robots in a CMDP formulation. However, their framework required cost coefficient engineering because they handled several constraints as rewards with appropriately set coefficients. Furthermore, their learning framework required considerable reward engineering due to multiple reward terms and a reference trajectory generator, whereas our framework requires only three reward terms and no demonstration data. Our comparison experiment with [10] also demonstrates that the controller obtained with our method is superior in the perspective of safety and task performance.
@@ IV-B Policy optimization
In the original paper [43], the IPO objective (10) is optimized with Proximal Policy Optimization (PPO) [49]. However, we utilized Trust Region Policy Optimization (TRPO) [48] because the optimizer’s explicit line search allows for more stable policy improvement and constraint feasibility checking  [^1] We also experimented with the PPO[49] optimizer and found several algorithmic advantages, such as ease of implementation and the availability to utilize recurrent neural networks, albeit with the trade-off of training stability.. The advantage functions $A^{\pi}$ and $A^{\pi}_{C_{k}}$ are computed using Generalized Advantage Estimation (GAE) [50] and neural network value functions. Each advantage function then undergoes additional standardization and zero-mean normalization before computing the policy gradients.
@@ VI Experimental Results
  • Results are after training with 1.6 billion time steps. Constraint thresholds are reported in the top row or inside the bracket. Bolded texts indicate violated or barely satisfied constraints.
@@ VI-A Implementation detail
The locomotion control policy runs at 100Hz. We deployed the trained controller on two real-world quadruped robots: Raibo [9] and Mini-cheetah [57]. For the deployment, the student policy network was reimplemented with the Eigen library in C++ and forward-passed in the Central Processing Unit (CPU). We recommend readers check the original paper [9, 57] for more details about the robot hardware.
@@ VI-B Framework Evaluation
The trained controllers were deployed on two real-world quadruped robots available in the lab: Raibo and Mini-cheetah. The controllers were extensively tested in various harsh terrains both indoor and outdoor to verify the robust control performance as shown in the previous work [6]. The testing environments included slippery hills, steep slopes, deformable terrains, stairs, a pile of leaves with hidden impediments, a moving cart, unstably piled planks, and even more. No matter how difficult the terrains were, Raibo and Mini-cheetah were able to walk steadily by modifying base motions and foot placements based solely on proprioceptive sensor data, as illustrated in Figure 4. Furthermore, Raibo could sprint over grassy terrain at 2m/s and climb slippery and steep slopes with an angle of around 32 at 1m/s which is roughly double the speed reported in the prior work [6].
@@ VI-D Comparison with the Reward-only Framework
We compare the generalizability of our framework and the reward-only framework, which is dominantly used in previous works on legged robot locomotion [6, 21, 8, 9]. The reward terms and hyperparameters used for the reward-only framework are based on the previous works by Ji et al.[8] and Choi et al. [9]. There were a total of twelve reward terms and they were designed for task completion (i.e., command tracking), physical constraints’ satisfaction (i.e., undesirable body contact, joint velocity, joint acceleration, joint torque, action smoothness), and style constraints’ satisfaction (i.e., air time, foot slippage, foot clearance, nominal configuration, hip joint usage, orthogonal velocity). Adequate reward coefficients for each term were laboriously engineered for robust control performance on the Raibo robot while satisfying all physical constraints. TRPO was used for policy optimization.
@@
Figure 7: Comparison of each reward component’s distribution with the reward-only framework. Red blocks show reward types with highly different reward distributions.
@@ VI-E2 Multi-head cost value function
  • Percentage in the ”Computation time” indicates the ratio of time spent on cost-critic versus others (e.g., simulation), where others are identical for all three methods.
The results are shown in Figure 8-B and Table VI. The multi-head method requires fewer trainable parameters than the single-head method. The multi-head cost value function requires 0.1 million trainable parameters, whereas single-head cost value functions require 1 million. Despite having considerably fewer neural network parameters to optimize, the policy optimized with the multi-head version yielded a similar reward as the single-head version (Figure 8-B). This indicates that sharing the neural network backbone is permissible for each cost-value function, and the network capacity is sufficient to model them. Because only the output layer size increases in proportion to the number of constraints, the suggested technique is highly scalable and effective for solving problems with numerous constraints.
@@ VII-A Comparison of velocity command tracking performance with the reward-only framework
Ours + L2 (Raibo) 0.498 ($\pm$ 0.308) 24.769 ($\pm$ 5.811)
Ours + Exp (Raibo) 0.367 ($\pm$ 0.290) 29.489 ($\pm$ 8.311)
Ours + Exp (Anymal C) 0.547 ($\pm$ 0.380) 60.310 ($\pm$ 14.978)
Reward-only (Raibo) 0.359 ($\pm$ 0.354) 30.773 ($\pm$ 10.181)
Figure 9: Maximum velocity command tracking results
As shown in Table VII, we can achieve the same command tracking performance with similar torque usage within our framework by applying the same exponential function for the command tracking reward and adjusting the reward coefficients accordingly ($k_{c}=200$, $\hat{k}_{\tau}=0.03$). When compared to a quadratic function, an exponential function has a steeper gradient near zero, resulting in more sensitive reward feedback for different command tracking errors. As a result, the robot can be trained to exhibit smaller command tracking errors. Additionally, similar maximum speeds can be achieved when given the maximum forward velocity (-1 $\mathrm{m}\mathrm{/}\mathrm{s}$ $\sim$ 2 $\mathrm{m}\mathrm{/}\mathrm{s}$) and turning rate (-1 $\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$ $\sim$ 1 $\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$) commands (Figure 9).
To verify the robot generalization ability with different reward designs, we re-trained the controller for the Anymal C robot with the same reward coefficients used for the Raibo robot but with an exponential reward kernel (Table VII). Similar results with Figure 6-A were obtained, indicating the maintenance of the robot generalization capability. This is because the critical component for achieving robot generalizability is the incorporation of constraints rather than specific reward designs. The proposed method offers strong robot generalizability as long as constraints are introduced and the effect is regardless of reward designs.
@@ VII-B Extra demonstration of constraint’s engineering advantages
  • ”(A)” indicates the engineer’s action. ”(R)” indicates the engineer’s response after reviewing the trained results for the action taken in ”(A)”.